# Adapting Physics-Informed Neural Networks To Optimize ODEs in Mosquito Population Dynamics

### v1

@@ 4.1 Forward Problem with T=2
In summary, $\text{OdePINN}_{\text{grad+causal}}$ successfully integrates both methodologies, ensuring adherence to the initial condition and minimizing ODE loss. This dual approach enables the model to more accurately approximate the true solution. These observations are further illustrated in Figure 2. Note that evaluation losses are computed across the entire period, with the weighting coefficients $\lambda_{i}$ maintained at 1.0; $\text{OdePINN}_{\text{orig}}$ are not shown as the losses are at different scales and its approximation is similar to $\text{OdePINN}_{\text{causal}}$. Figure 2(a) demonstrates that the $\text{OdePINN}_{\text{grad+causal}}$ model achieves a significantly low final total loss of $1.6\cdot 10^{-4}$, which is considerably less than that of the $\text{OdePINN}_{\text{causal}}$ at $2.2\cdot 10^{-3}$. The remaining models display higher losses of 0.046 and 0.056, with the baseline model showing slightly better performance. Comparison of errors across the models indicates that Gradient Balancing substantially enhances convergence towards the data loss term $\mathcal{L}_{data}$. The data losses for $\text{OdePINN}_{\text{grad}}$ and $\text{OdePINN}_{\text{grad+causal}}$ are recorded at $1.4\cdot 10^{-8}$ and $8.9\cdot 10^{-9}$, respectively, which are significantly lower than the approximately $10^{-2}$ or $10^{-3}$ observed in the other models. For an in-depth analysis of the method, see [36].
@@
However, the Gradient Balancing technique on its own does not address the issue of causal effects: Causal Training is essential in this case. Gradient Balancing averages gradients from different times, leading to scenarios like those in Figure 2(b), where at later time $t=1.5$ or $t=2.0$, the model can quickly minimize the physics constraints to a negligible level of $10^{-5}$ at an early stage in the training. Notably, within the domain $t\in[1.0,2.0]$, the model converges to the null solution (as illustrated in Figure 1), which satisfies the ODE system. However, at earlier times, such as $t=0$ or $t=0.5$, the model consistently struggles throughout the training duration. We hypothesize that this premature convergence of the ODE constraints at later times traps the model in a local minimum, preventing it from satisfying the constraints at earlier times and thus resulting in an inaccurate solution approximation. To address this, we employ a Causal Training strategy that trains the physics-laws term using residual points drawn from a progressively expanding interval, starting from the 1,000th to the 200,000th step. This growing interval strategy ensures that the model adheres to the system dynamics at earlier times before progressing to later times, thereby respecting the causal effects. The domain in which the model complies with the differential equations broadens as the training advances, as depicted in Figure 2(c). Outside this domain, the model’s behavior remains arbitrary. Ultimately, the $\text{OdePINN}_{\text{grad+causal}}$ model effectively combines these techniques, effectively minimizing both data and physics losses to achieve a highly accurate solution approximation.
4.2 Forward Problem $T=20$
@@
Figure 3: Lorenz ODE system, forward problem, $U$ approximation of the system state.
@@
Figure 3 illustrates the solution approximated by the proposed framework, accompanied by plots of the training losses and Root Mean Squared Error (RMSE) relative to the ground truth data. The model demonstrates a fairly good prediction of the system’s evolution. Initially, the RMSE is approximately $5\cdot 10^{-4}$, but it exponentially increases as $t$ progresses, where the RMSE escalates to $10^{-3}$ at $t=8$, to $10^{-1}$ at $t=13$, and reaches an error magnitude of $10^{1}$ by the end of the period. The final four subdomains, as depicted in figure 3 (12.5 to 20.0), show notable approximation errors. This substantial error accumulation towards the end of the period is understandable given that the Lorenz system is highly sensitive to initial conditions where minor predictive inaccuracies can significantly deviate the future states. As a result, initial training errors rapidly accumulate over time, culminating in an RMSE of up to 10 by the end of the time period. The training data loss remains below $10^{-4}$ throughout, with many instances dropping to the $10^{-5}$ level. Furthermore, the errors related to physical constraints are consistently maintained at the $10^{-3}$ level, indicating that the physical laws are satisfied with ODE loss approaching zero. Notably, the losses peak in regions of the solution characterized by sharp changes, consequently resulting in a steep increase in RMSE. On a positive front, the framework exhibits consistent performance across all subdomains. However, the clear accumulation of errors highlights a potential limitation of the technique, suggesting areas for further improvement.
4.3 Inverse Problem
(a) Ground truth $u$ and data provided to PINN.
(b) $\Theta$ approximation of the physics parameters $(\sigma,\rho,\beta)$.
Figure 4: OdePINN framework solves the Lorenz system inverse problem.
@@ 4.3 Inverse Problem
In this section, we illustrate the application of the proposed framework to an inverse problem scenario, using significantly more data. Here, the framework is employed to predict the physics parameters and simultaneously interpolating the system state from a limited number of observations of the system’s state. We conduct these experiments using the Lorenz system over the time domain $t\in[0,2.0]$, with time-varying physical parameters defined as $\sigma=\frac{10}{2}\sin\left(2\pi t\right)+10$, $\rho=\frac{28}{5}\sin\left(2\pi t+\frac{\pi}{2}\right)+28$, and $\beta=\frac{8}{3}$. These formulas are unknown to and to be learnt by the framework. The initial conditions are set to $(1,1,1)$, consistent with previous experiments. The dataset for training comprises 21 simulated data points that are evenly distributed across the input domain, resulting in a dataset dimension of $(21,3)$. Figure 4(a) presents both the reference solution discussed earlier [78] and the simulated data points.
@@
Model Metric $x$ $y$ $z$ $u$ $\sigma$ $\rho$ $\beta$ $\theta$
orig RMSE 4.4453 4.9325 1.6939 3.9564 5.4294 90.5381 3.8850 52.4141
baseline 3.4995 2.8182 0.6591 2.6219 8.4507 41.5828 1.3683 24.5114
@@
causal 0.1363 0.0949 0.0692 0.1039 1.3853 2.4592 0.0740 1.6302
[…]
orig MAE 1.7002 1.0028 0.3667 1.0232 3.6059 47.0012 1.4000 17.3357
baseline 0.9579 0.6687 0.2430 0.6232 3.1751 10.9855 0.5066 4.8891
grad 0.0207 0.0759 0.0449 0.0472 0.2116 0.2078 0.0611 0.1602
causal 0.0844 0.0654 0.0452 0.0650 0.7099 1.3002 0.0472 0.6858
[…]
orig MDAPE 0.0063 0.0064 0.0015 0.0025 0.1271 0.0127 0.0058 0.0200
baseline 0.0019 0.0029 0.0016 0.0019 0.0382 0.0114 0.0086 0.0162
grad 0.0001 0.0002 0.0001 0.0001 0.0005 0.0005 0.0003 0.0004
causal 0.0004 0.0012 0.0016 0.0013 0.0175 0.0044 0.0074 0.0090
[…]
orig nRMSE 0.2562 0.2450 0.0703 0.2086 2.7147 16.1675 3.8850 9.7271
baseline 0.2017 0.1400 0.0274 0.1426 4.2253 7.4255 1.3683 4.9955
[…]
causal 0.0021 0.0025 0.0008 0.0019 0.2210 0.0608 0.0156 0.1326
grad+causal 0.0079 0.0162 0.0140 0.0132 0.5519 0.1572 0.3124 0.3772
Table 1: Approximation errors in the inverse problem with Lorenz system where smallest error values are best. Bold text highlight the best performing models with respect to a specific metric and Underlined numbers represent the second best performing model.
Figure 4(b) (in the appendix) displays the dynamical system’s parameters as approximated by the neural network $\Theta$. While all models demonstrate an ability to capture the general trends of the parameters, they also exhibit inherent inaccuracies. Notably, the original and baseline models show substantial errors particularly in regions where the derivatives of $u$ with respect to time $t$ are remarkably high, as can be seen within the interval $[0,0.5]$ or by $t=2.0$ in figure 4(a). Among the models, $\text{OdePINN}_{\text{grad}}$ achieves the most accurate parameter approximations throughout most of the domain. However, it encounters difficulties at the period’s end, near $t=2.0$. In this inverse setting, the system’s parameters are not predefined and are subject to variation, leading to non-unique solutions for $u$ and $\theta$ in the absence of sufficient data. Consequently, $\text{OdePINN}_{\text{grad}}$ tends to favor solutions with smaller magnitude derivatives, which can result in significant discrepancies when compared to the reference solution. Conversely, $\text{OdePINN}_{\text{causal}}$ typically exhibits error accumulation towards higher values of $t$, with noticeable inaccuracies in estimating parameters such as $\sigma$ and $\rho$ beyond $t=1.15$, as illustrated in figure 4(b). This pattern of error propagation is a recurring observation throughout our study. The combined model, $\text{OdePINN}_{\text{grad+causal}}$, mirrors the performance of $\text{OdePINN}_{\text{grad}}$, maintaining reasonable accuracy until the domain’s far end. The approximation $U$ by the models can be found in figure 4(a) in the Appendix B.
Table 1 presents a detailed analysis of the errors associated with both the interpolation of $u$ and the estimation of parameters $\theta$, benchmarked against the ground truth data, errors calculated in metrics presented in Appendix A. The original model registers highest errors, with RMSE values of 3.9564 for $u$ and 52.4141 for $\theta$, alongside MDAPE of 2.6219 and 24.5114, respectively. Compared to the original model, the baseline model shows slight improvements but still incurs relatively high errors. Conversely, $\text{OdePINN}_{\text{grad}}$ obtained significant improvements in accuracy, recording the lowest errors across all evaluated metrics for both $u$ and $\theta$. It achieves an RMSE of 0.2525 for $u$ and an impressively low 0.7195 for $\theta$. Additionally, it reports MAE values of 0.0472 and 0.1602, and MDAPE scores below 0.01% for both variables. The causal model records the best RMSE for $u$ at 0.1039, although its performance on other metrics does not reach the levels achieved by the grad model. $\text{OdePINN}_{\text{grad+causal}}$, which integrates the approaches of the preceding two models, does not achieve top scores in any specific category but secures the second-best results, including impressive MDAPE values below 0.01% and 0.06%. This final training stage of $\text{OdePINN}_{\text{grad+causal}}$ mirrors the training process of the entire $\text{OdePINN}_{\text{grad+causal}}$ model, albeit with an improved initialization and a reduced number of training steps.
@@
5 Case Study-Based Validation
In addition to a theoretical validation, our PINN framework is also validated in the practical environment of dynamical modelling of the mosquito population. Here, our approach is adopted on the ODE-based model of mosquito population dynamics proposed in [80]. The model divides the mosquito life cycle into 10 stages: Egg ($E$), Larva ($L$), Pupa ($P$), Emerging Adults ($A_{em}$), Nulliparous Bloodseeking Adults ($A_{b1}$), Nulliparous Gestating Adults ($A_{g1}$), Nulliparous Ovipositing Adults ($A_{o1}$), Parous Bloodseeking Adults ($A_{b2}$), Parous Gestating Adults ($A_{g2}$) and Parous Ovipositing Adults ($A_{o2}$). The 10 stages are related via ordinary equations as show in equation 31, with parameters explained in table 4 in Appendix C. System parameters are modelled as a function depending on temperature with data acquired for Culex pipiens spieces from [11].
@@ 5 Case Study-Based Validation
5.1 Forward Problem
Figure 5: Mosquito ODE system, forward problem, $U$ approximation of the solution.
For the forward problem, the goal was to solve the mosquito population dynamics using only a single data point at the beginning, known as the initial condition. Specifically, we employ a numerical method [78] for solving ordinary differential equations (ODEs), available in Python Scipy library [79] , to simulate data over three years ($t$ ranging from 00 to $1096$ days). We use the data point at $t=730$ (two years into the simulation) as the starting point and trained a PINN for the period from $730$ to $1,096$ days. The simulated data within this period serves as the ground truth solution against which we evaluate the model’s performance. The advanced techniques presented in the section 3, including ODE normalization, domain decomposition, gradient balancing and causal training, were implemented.
@@ 5.1 Forward Problem
Table 2: Approximation Errors for Mosquito ODE Solution.
The results for solving the mosquito population dynamics as a forward problem are shown in figure 5, with error measurements provided in Table 2. The trained neural network shows impressive performance in extrapolating and providing a fairly accurate approximation of the solution in figure 5. However, towards the end of the time period, accumulated error can be seen in Ag1 and Ag2 stages of mosquitoes development.
@@
5.2 Inverse Problem
In solving inverse problems, the goal was to predict 10 specific parameters identified in table 4 using available data. Three parameters ($\gamma_{Aem}$, $\gamma_{Ab}$ and $\gamma_{Ao}$) are treated as constants and are represented by learnable parameters that remain constant over time. The remaining seven parameters ($f_{E}$, $f_{P}$, $f_{L}$, $f_{Ag}$, $m_{L}$, $m_{P}$ and $m_{A}$) vary over time with neural networks using time as input to estimate these parameters. This experiment was conducted under the same conditions and temperatures as previous experiments but here, it was necessary to generate simulation data over a three-year period, from day 0 to day 1096, using the numerical method in [78]. For the data condition loss, we used daily data from $t=730$ to $t=1096$, totaling 367 days times 10 data points. Both the solution $u$ and the ODE parameters $\theta$ were obtained from the simulation data as ground truth.
@@ 5.2 Inverse Problem
Figure 6: Mosquito ODE system, inverse problem, $\Theta$ approximation of the system’s parameters.
@@
Table 3: Mosquito ODE System’s Parameter Approximation Errors
Figure 6 shows the system parameters learnt from the PINN framework. The plots generally illustrate accurate parameter approximations although there are noticeable inaccuracies, particularly for parameters $m_{L}$ and $m_{P}$ within the domain $[943,1064]$. The error metrics detailed in table 3, reveal an average RMSE of 0.132615, with $m_{L}$ and $m_{P}$ contributing the highest errors at 0.286959 and 0.291548, respectively. The lowest RMSE values are for $f_{Ag}$ and $f_{L}$ at 0.002769 and 0.002816, respectively. The MAE closely follows the RMSE trends, averaging at 0.044757. The highest MAE values correspond to $m_{L}$ and $m_{P}$ (0.137099 and 0.139703), and the lowest to $f_{Ag}$ and $f_{L}$ (0.001982 and 0.001784). The MDAPE averages at 5.9%, with $m_{A}$ showing the lowest error at 3.3566% and the three constants ($\gamma_{Aem}$, $\gamma_{Ab}$, and $\gamma_{Ao}$) ranging between 3.62% and 3.87%. The errors for $m_{L}$ and $m_{P}$ are notably higher at 12.59% and 25.91%, respectively.
@@
We interpret these results as follows: in colder temperatures, within the day range of 943 to 1064, where calculated air temperature is below $5.0^{\circ}C$, the mosquito population decreases rapidly to zero (with the exception of $Ag1$ and $Ag2$) as can be seen in figure 5. This situation makes it challenging to gather useful information, negatively affecting the ability to determine several system parameters, especially $m_{L}$ and $m_{P}$ as these parameters are primarily derived from equations related to $\frac{dL}{dt}$ and $\frac{dP}{dt}$. Conversely, parameters $\gamma_{Aem}$, $\gamma_{Ab}$ and $\gamma_{Ao}$ remain constant over time, which makes them simpler to accurately estimate, as seen in their relatively low MDAPE. Meanwhile, parameters like $m_{A}$, $f_{Ag}$ which are involved in several differential equations, tend to receive more consistent and stable information and gradients. Mathematically, the target parameters are not structurally identifiable across the entire input domain due to the system setup (see Appendix D), which can result in inaccuracies in the approximations made by PINNs.
6 Discussion and Future Work
@@
In this work, we introduced a hybrid framework based on Physics informed Neural Networks that integrates physical laws into data-driven machine learning models. This particular framework is designed to manage systems of ordinary differential equations, validated using a case study modeling mosquito population dynamics. The approach employs a multi-task learning strategy, incorporating multiple components in the objective function: one for data fitting, several for weakly enforcing physical knowledge, with separate functions for the remaining conditions. The framework includes several advanced techniques such as domain decomposition, ODE normalization, gradient balancing, and causal training. We evaluated the effectiveness of our approach through an ablation study using the Lorenz system before introducing the problem of a model for the dynamics of mosquito populations.
@@
Our findings demonstrate that ODE Normalization and Gradient Balancing played an important role in stabilizing the training process. These techniques ensured that no individual component of the loss function disproportionately influenced the optimization, thus preventing premature convergence to suboptimal solutions. Causal Training preserves the temporal causality inherent in the dynamical system, critical for achieving accurate model predictions, especially in scenarios where extrapolation beyond the scope of the training data is necessary. Domain Decomposition effectively manages significantly large input domains, particularly in forward problem scenarios. The results also confirm the framework’s efficacy in modeling mosquito population dynamics, highlighting its potential for application within the field of ecology.

### v2

@@ Methods
In summary, $\text{OdePINN}_{\text{grad+causal}}$ successfully integrates both methodologies, ensuring adherence to the initial condition and minimizing ODE loss. This dual approach enables the model to more accurately approximate the true solution. These observations are further illustrated in Fig 2. Note that evaluation losses are computed across the entire period, with the weighting coefficients $\lambda_{i}$ maintained at 1.0; $\text{OdePINN}_{\text{orig}}$ are not shown as the losses are at different scales and its approximation is similar to $\text{OdePINN}_{\text{causal}}$. Fig 2a demonstrates that the $\text{OdePINN}_{\text{grad+causal}}$ model achieves a significantly low final total loss of $1.6\cdot 10^{-4}$, which is considerably less than that of the $\text{OdePINN}_{\text{causal}}$ at $2.2\cdot 10^{-3}$. The remaining models display higher losses of 0.046 and 0.056, with the baseline model showing slightly better performance. Comparison of errors across the models indicates that Gradient Balancing substantially enhances convergence towards the data loss term $\mathcal{L}_{data}$. The data losses for $\text{OdePINN}_{\text{grad}}$ and $\text{OdePINN}_{\text{grad+causal}}$ are recorded at $1.4\cdot 10^{-8}$ and $8.9\cdot 10^{-9}$, respectively, which are significantly lower than the approximately $10^{-2}$ or $10^{-3}$ observed in the other models. Despite achieving a significantly lower data loss, the total loss of $\text{OdePINN}_{\text{grad}}$ is $6.3\cdot 10^{-2}$, slightly higher than the baseline model’s $4.5\cdot 10^{-2}$, due to a higher ODE loss. This technique enables the neural network to better fit the initial condition but simultaneously makes it harder to satisfy the ODE loss, resulting in higher ODE and overall losses. Greater detail on the usage and analysis of this method is provided in [36]. Ultimately, the root mean squared error between the grad model and the true solution is 11.9, slightly better than the baseline’s 12.18, though both remain far from a perfect solution.
Fig 2: Loss analysis of OdePINN framework with Lorenz system, $t\in[0,2.0]$. (a) Loss Terms across final models selecting through Early Stopping; the bars are presented in a logscale pointing downwards, with the lower the value the better the performance. (b) ODE errors of the $\text{OdePINN}_{\text{grad}}$ model at different time $t$ during the training, motivating the need for Causal Training; (c) The $x$-value Approximation Solution $U$ of the model $\text{OdePINN}_{\text{grad+causal}}$ at different steps during the training.
@@
However, the Gradient Balancing technique on its own does not address the issue of causal effects: Causal Training is essential in this case. Gradient Balancing averages gradients from different times, leading to scenarios in Fig 2b, where at later time $t=1.5$ or $t=2.0$, the model can quickly minimize the physics constraints to a negligible level of $10^{-5}$ at an early stage in the training. Notably, within the domain $t\in[1.0,2.0]$, the model converges to the null solution (as illustrated in Fig 1), which satisfies the ODE system. However, at earlier times, such as $t=0$ or $t=0.5$, the model consistently struggles throughout the training duration. We hypothesize that this premature convergence of the ODE constraints at later times traps the model in a local minimum, preventing it from satisfying the constraints at earlier times and thus resulting in an inaccurate solution approximation. To address this, we employ a Causal Training strategy that trains the physics-laws term using residual points drawn from a progressively expanding interval, starting from the 1,000th to the 200,000th step. This growing interval strategy ensures that the model adheres to the system dynamics at earlier times before progressing to later times, thereby respecting the causal effects. The domain in which the model complies with the differential equations broadens as the training advances, as depicted in Fig 2c. Outside this domain, the model’s behavior remains arbitrary. Ultimately, the $\text{OdePINN}_{\text{grad+causal}}$ model effectively combines these techniques, effectively minimizing both data and physics losses to achieve a highly accurate solution approximation.
Forward Problem $T=20$
@@
Fig 3 displays the the results when the domain is larger where none of the models achieve a close approximation to the reference solution. The two models $\text{OdePINN}_{\text{baseline}}$ and $\text{OdePINN}_{\text{grad}}$ demonstrate some level of ability to learn the initial condition and marginally adhere to the ODE equations. However, their errors grow sufficiently large by $t=0.75$, leading them to diverge significantly from the true solution, eventually converging to a trivial constant solution beyond $t=3$. Other models, $\text{OdePINN}_{\text{orig}}$, $\text{OdePINN}_{\text{causal}}$ and $\text{OdePINN}_{\text{grad + causal}}$ exhibit a similar pattern as observed in the $T=2$ experiment, converging to the null solution. The models’ inaccuracies can be attributed to the increased domain size, which raises the solution’s complexity. Therefore, it was necessary to apply Domain Decomposition to manage the larger domain size.
Fig 3: Lorenz ODE system, forward problem, $U$ approximation of the system state from the first 5 models, using the first five models, excluding the Domain Decomposition model.
@@
Fig 4 illustrates the solution approximated by the proposed framework, accompanied by plots of the training losses and Root Mean Squared Error (RMSE) relative to the ground truth data. The model demonstrates a fairly good prediction of the system’s evolution. Initially, the RMSE is approximately $5\cdot 10^{-4}$, but it exponentially increases as $t$ progresses, where the RMSE escalates to $10^{-3}$ at $t=8$, to $10^{-1}$ at $t=13$, and reaches an error magnitude of $10^{1}$ by the end of the period. The final four subdomains, as depicted in Fig 4 (12.5 to 20.0), show notable approximation errors. This substantial error accumulation towards the end of the period is understandable given that the Lorenz system is highly sensitive to initial conditions where minor predictive inaccuracies can significantly deviate the future states. As a result, initial training errors rapidly accumulate over time, culminating in an RMSE of up to 10 by the end of the time period. The training data loss remains below $10^{-4}$ throughout, with many instances dropping to the $10^{-5}$ level. Furthermore, the errors related to physical constraints are consistently maintained at the $10^{-3}$ level, indicating that the physical laws are satisfied with ODE loss approaching zero. Notably, the losses peak in regions of the solution characterized by sharp changes, consequently resulting in a steep increase in RMSE. On a positive front, the framework exhibits consistent performance across all subdomains. However, the clear accumulation of errors highlights a potential limitation of the technique, suggesting areas for further improvement.
Fig 4: Lorenz ODE system, forward problem, $U$ approximation of the system state, using the domain composition model.
Fig 5 plots the relationship between the number of training steps and the RMSE across a range of subdomains. The training steps increase linearly with the number of subdomains as each subdomain requires approximately 200,000 to 300,000 training steps, equivalent to around 30 minutes on the GPU NVIDIA GeForce RTX 4090 hardware used for all experiments. The figure demonstrates a clear trade-off between the number of subdomains and model accuracy: as the number of subdomains rise, the RMSE decreases exponentially, but at the cost of longer training times. When the number of subdomains is low, such as 1 or 5, the training domain remains too large, causing the model to converge to the null solution, yielding an RMSE of 13.4. With an increase in subdomains, starting from 10, the training domain becomes smaller, reducing the solution complexity within each subdomain. This reduction leads to a RMSE decrease, reaching 4.1 with 10 subdomains and 1.13 with 40 subdomains. These results indicate that cumulative errors can be mitigated by increasing the number of subdomains, allowing the model to focus on improving accuracy at finer levels of detail.
Fig 5: Trade-off between the number of subdomains and accuracy.
Inverse Problem
@@
In this section, we illustrate the application of the proposed framework to an inverse problem scenario, using significantly more data. Here, the framework is employed to predict the physics parameters and simultaneously interpolating the system state from a limited number of observations of the system’s state. We conduct these experiments using the Lorenz system over the time domain $t\in[0,2.0]$, with time-varying physical parameters defined as $\sigma=\frac{10}{2}\sin\left(2\pi t\right)+10$, $\rho=\frac{28}{5}\sin\left(2\pi t+\frac{\pi}{2}\right)+28$, and $\beta=\frac{8}{3}$. These formulas are unknown to and are to be learnt by the framework. The initial conditions are set to $(1,1,1)$, consistent with previous experiments. The dataset for training comprises 21 simulated data points that are evenly distributed across the input domain, resulting in a dataset dimension of $(21,3)$. Fig 6a presents both the reference solution discussed earlier [78] and the simulated data points.
Fig 6: OdePINN framework solves the Lorenz system inverse problem. (a): Ground truth $u$ and data provided to PINN; (b): $\Theta$ approximation of the physics parameters $(\sigma,\rho,\beta)$.
@@
Fig 6b displays the dynamical system’s parameters as approximated by the neural network $\Theta$. While all models demonstrate an ability to capture the general trends of the parameters, they also exhibit inherent inaccuracies. Notably, the original and baseline models show substantial errors particularly in regions where the derivatives of $u$ with respect to time $t$ are remarkably high, as can be seen within the interval $[0,0.5]$ or by $t=2.0$ in Fig 6a. Among the models, $\text{OdePINN}_{\text{grad}}$ achieves the most accurate parameter approximations throughout most of the domain. However, it encounters difficulties at the period’s end, near $t=2.0$. In this inverse setting, the system’s parameters are not predefined and are subject to variation, leading to non-unique solutions for $u$ and $\theta$ in the absence of sufficient data. Consequently, $\text{OdePINN}_{\text{grad}}$ tends to favor solutions with smaller magnitude derivatives, which can result in significant discrepancies when compared to the reference solution. Conversely, $\text{OdePINN}_{\text{causal}}$ typically exhibits error accumulation towards higher values of $t$, with noticeable inaccuracies in estimating parameters such as $\sigma$ and $\rho$ beyond $t=1.15$, as illustrated in Fig 6b. This pattern of error propagation is a recurring observation throughout our study. The combined model, $\text{OdePINN}_{\text{grad+causal}}$, mirrors the performance of $\text{OdePINN}_{\text{grad}}$, maintaining reasonable accuracy until the domain’s far end. The approximation $U$ by the models can be found in S1 Fig..
Table 1 presents a detailed analysis of the errors associated with both the interpolation of $u$ and the estimation of parameters $\theta$, benchmarked against the ground truth data, errors calculated in metrics presented in S1 Appendix.. The original model registers highest errors, with RMSE values of 3.9564 for $u$ and 52.4141 for $\theta$, alongside MDAPE of 2.6219 and 24.5114, respectively. Compared to the original model, the baseline model shows slight improvements but still incurs relatively high errors. Conversely, $\text{OdePINN}_{\text{grad}}$ obtained significant improvements in accuracy, recording the lowest errors across all evaluated metrics for both $u$ and $\theta$. It achieves an RMSE of 0.2525 for $u$ and an impressively low 0.7195 for $\theta$. Additionally, it reports MAE values of 0.0472 and 0.1602, and MDAPE scores below 0.01% for both variables. The causal model records the best RMSE for $u$ at 0.1039, although its performance on other metrics does not reach the levels achieved by the grad model. $\text{OdePINN}_{\text{grad+causal}}$, which integrates the approaches of the preceding two models, does not achieve top scores in any specific category but secures the second-best results, including impressive MDAPE values below 0.01% and 0.06%. This final training stage of $\text{OdePINN}_{\text{grad+causal}}$ mirrors the training process of the entire $\text{OdePINN}_{\text{grad+causal}}$ model, albeit with an improved initialization and a reduced number of training steps.
{adjustwidth}
@@
Table 1: Approximation errors in the inverse problem with Lorenz system where smallest error values are best.
Bold text highlight the best performing models with respect to a specific metric and Underlined numbers represent the second best performing model.
@@
Case Study-Based Validation
In addition to a theoretical validation, our PINN framework is also validated in the practical environment of dynamical modelling of the mosquito population. Here, our approach is adopted on the ODE-based model of mosquito population dynamics proposed in [80]. The model divides the mosquito life cycle into 10 stages: Egg ($E$), Larva ($L$), Pupa ($P$), Emerging Adults ($A_{em}$), Nulliparous Bloodseeking Adults ($A_{b1}$), Nulliparous Gestating Adults ($A_{g1}$), Nulliparous Ovipositing Adults ($A_{o1}$), Parous Bloodseeking Adults ($A_{b2}$), Parous Gestating Adults ($A_{g2}$) and Parous Ovipositing Adults ($A_{o2}$). The 10 stages are related via ordinary equations as show in Eq (31), with parameters explained in S1 Table.. A diagram illustrating the stages and transitions is available in S2 Fig.. System parameters are modelled as a function depending on temperature with data acquired for Culex pipiens spieces from [11].
@@
Forward Problem
For the forward problem, the goal was to solve the mosquito population dynamics using only a single data point at the beginning, known as the initial condition. Specifically, we employ a numerical method [78] for solving ordinary differential equations (ODEs), available in Python Scipy library [79] , to simulate data over three years ($t$ ranging from 00 to $1096$ days). We use the data point at $t=730$ (two years into the simulation) as the starting point and trained a PINN for the period from $730$ to $1,096$ days. The simulated data within this period serves as the ground truth solution against which we evaluate the model’s performance. The advanced techniques presented in the section Methods, including ODE normalization, domain decomposition, gradient balancing and causal training, were implemented.
@@
The results for solving the mosquito population dynamics as a forward problem are shown in Fig 7, with error measurements provided in Table 2. The trained neural network shows impressive performance in extrapolating and providing a fairly accurate approximation of the solution in Fig 7. However, towards the end of the time period, accumulated error can be seen in Ag1 and Ag2 stages of mosquitoes development.
Fig 7: Mosquito ODE system, forward problem, $U$ approximation of the solution.
Table 2: Approximation Errors for Mosquito ODE Solution.
@@
Inverse Problem
In solving inverse problems, the goal was to predict 10 specific parameters using available data. Three parameters ($\gamma_{Aem}$, $\gamma_{Ab}$ and $\gamma_{Ao}$) are treated as constants and are represented by learnable parameters that remain constant over time. The remaining seven parameters ($f_{E}$, $f_{P}$, $f_{L}$, $f_{Ag}$, $m_{L}$, $m_{P}$ and $m_{A}$) vary over time with neural networks using time as input to estimate these parameters. This experiment was conducted under the same conditions and temperatures as previous experiments but here, it was necessary to generate simulation data over a three-year period, from day 0 to day 1096, using the numerical method in [78]. For the data condition loss, we used daily data from $t=730$ to $t=1096$, totaling 367 days times 10 data points. Both the solution $u$ and the ODE parameters $\theta$ were obtained from the simulation data as ground truth.
@@
Fig 8 shows the system parameters learnt from the PINN framework. The plots generally illustrate accurate parameter approximations although there are noticeable inaccuracies, particularly for parameters $m_{L}$ and $m_{P}$ within the domain $[943,1064]$. The error metrics detailed in Table 3, reveal an average RMSE of 0.132615, with $m_{L}$ and $m_{P}$ contributing the highest errors at 0.286959 and 0.291548, respectively. The lowest RMSE values are for $f_{Ag}$ and $f_{L}$ at 0.002769 and 0.002816, respectively. The MAE closely follows the RMSE trends, averaging at 0.044757. The highest MAE values correspond to $m_{L}$ and $m_{P}$ (0.137099 and 0.139703), and the lowest to $f_{Ag}$ and $f_{L}$ (0.001982 and 0.001784). The MDAPE averages at 5.9%, with $m_{A}$ showing the lowest error at 3.3566% and the three constants ($\gamma_{Aem}$, $\gamma_{Ab}$, and $\gamma_{Ao}$) ranging between 3.62% and 3.87%. The errors for $m_{L}$ and $m_{P}$ are notably higher at 12.59% and 25.91%, respectively.
Fig 8: Mosquito ODE system, inverse problem, $\Theta$ approximation of the system’s parameters.
Table 3: Mosquito ODE System’s Parameter Approximation Errors
@@
We interpret these results as follows: in colder temperatures, within the day range of 943 to 1064, where calculated air temperature is below $5.0^{\circ}C$, the mosquito population decreases rapidly to zero (with the exception of $Ag1$ and $Ag2$) as can be seen in Fig 7. This situation makes it challenging to gather useful information, negatively affecting the ability to determine several system parameters, especially $m_{L}$ and $m_{P}$ as these parameters are primarily derived from equations related to $\frac{dL}{dt}$ and $\frac{dP}{dt}$. Conversely, parameters $\gamma_{Aem}$, $\gamma_{Ab}$ and $\gamma_{Ao}$ remain constant over time, which makes them simpler to accurately estimate, as seen in their relatively low MDAPE. Meanwhile, parameters like $m_{A}$, $f_{Ag}$ which are involved in several differential equations, tend to receive more consistent and stable information and gradients. In fact, the target parameters are not identifiable across the entire input domain due to the system setup (see S3 Fig.), which can result in inaccuracies in the approximations made by PINNs.
Discussion and Future Work
@@
In this work, we introduced a hybrid framework based on Physics informed Neural Networks that integrates physical laws into data-driven machine learning models. This particular framework is designed to manage systems of ordinary differential equations, validated using a case study modelling mosquito population dynamics. The approach employs a multi-task learning strategy, incorporating multiple components in the objective function: one for data fitting, several for weakly enforcing physical knowledge, with separate functions for the remaining conditions. The framework includes several advanced techniques such as domain decomposition, ODE normalization, gradient balancing, and causal training. We evaluated the effectiveness of our approach through an ablation study using the Lorenz system before introducing the problem of a model for the dynamics of mosquito populations.
The application of PINNs to mosquito population dynamics represents a significant advancement in ecological modelling for a number of reasons. Traditional data-driven models often require extensive data and are limited to capturing the nonlinear, multi-scale behaviours seen in real-world mosquito populations. However, PINNs integrate domain-specific knowledge directly into the learning process by enforcing ODE constraints. This is crucial for mosquito population dynamics because it ensures that the model respects biological principles even with limited data.
@@
Our findings demonstrate that ODE Normalization and Gradient Balancing played an important role in stabilizing the training process. These techniques ensured that no individual component of the loss function disproportionately influenced the optimization, thus preventing premature convergence to suboptimal solutions. Causal Training preserves the temporal causality inherent in the dynamical system, critical for achieving accurate model predictions, especially in scenarios where extrapolation beyond the scope of the training data is necessary. Domain Decomposition effectively manages significantly large input domains, particularly in forward problem scenarios. The results also confirm the framework’s efficacy in modelling mosquito population dynamics, highlighting its potential for application to other ODE-based ecological models, which are commonly applied not only to mosquito populations but also to other vector species, such as ticks [81], also agricultural and forest pests, including Drosophila suzukii [82], Batrocera oleae [83] and bark beetles [84].